Order: 3,500 mL D5LR IV to run at 160 mL/h. Drop factor: 15 gtt/mL. What is the drip rate?
Answer #1

Rate = 160 ml/hour

Drop factor = 15 gtt/ml

So 1 ml= 15 gtt

160 ml= 160 * 15 = 2400 gtt

Drip rate = 2400 gtt per hour

= 2400/60 = 40 gtt/minute